# Enhancement of the molecular electron chirality by electronic excitation

### v1

@@ I Introduction
It is well-known that some molecules have chiral structures, while the electron also has its chirality, i.e. the handedness of the electron can be defined. A free electron does not have a specific charge distribution and the handedness of the electron cannot be defined by the distribution. However, the electron has the spin, and the inner product of the spin and momentum can define one type of handedness quantity of the electron. This definition of the handedness is the helicity. When the external field is negligible, this helicity is a conserved quantity in time, since the helicity commutes with the Hamiltonian. However, the helicity has a different ratio of the electron handedness in a different inertial frame and is not Lorentz invariant definition. The chirality of the electron is defined as the same handedness ration in any inertial frame. This is a relativistic extension of the helicity as explained in the next section.
@@
Nature chooses the chirality as the handedness quantity of electrons. First, the weak interaction, which is one of the four fundamental interactions in particle physics, distinguish left- and right-handed electrons by the chirality. For example, this interaction is mediated by W and Z gauge bosons, and the W boson has the coupling with left-handed electrons and not with right-handed ones. In another example of the distinction, the torque density for the electron spin in quantum field theory is predicted to be dependent on the electron chirality density [1, 2]. In this theoretical prediction, the local torque contribution, which is named the zeta force, is given as the gradient of the electron chirality density. Although this contribution of the torque density has not been observed in experiments, this is confirmed by numerical simulations [3].
@@
Although the physical property of the electron is different in its chirality, free electrons lose their polarization in chirality sooner or later due to the chirality flip by the electron mass. For example, if electrons are initially polarized more in left-handed chirality, it attenuates to no polarization state. The left-handed component flips to the right-handed one more than the inverse process (from right-handed to left-handed), since the ratio of the left-handed component is larger than the right-handed one.
@@
However, electrons in chiral molecules, which are parity-violating objects, are considered to be polarized in the chirality, since a state polarized in the electron chirality may be the most stable in the steady state due to the spin-orbit interaction. Even though the electron mass flips the electron chirality, the attenuation goal is the polarized state in the chirality.
@@
One phenomenon related to the electron chirality is the parity-violating energy difference (PVED) between enantiomers. This energy difference is induced by the interaction of Z-boson exchange between nuclei and electrons and is proportional to the electron chirality density difference at nuclei between enantiomers. Although this energy difference has not been confirmed in observational experiments despite hard efforts [4], theoretical works with quantum chemistry computations confirm nonzero PVED in many chiral molecules [5, 6, 7, 8]. Nevertheless, this energy difference is related to only the electron chirality at nuclei. The electron chirality for the entire molecule of H2Te2 is reported to be nonzero [9]. Then, in Ref. [6, 10, 11], it is systematically shown that chiral molecules have generally a nonzero integrated value of electron chirality density over the entire molecule. The existence of this chirality is named Electron Chirality in Chiral Molecules (ECCM) [11].
@@
The value of ECCM is not large, and phenomena related to ECCM are not so many. Nevertheless, some phenomena are studied by our group. One topic is the homochirality in nature [12]. In an enantiomeric pair, the number of the left-handed electron is different, though the total electron number is the same. Hence, one enantiomer has a larger reaction rate of weak interaction than the other one. As a result, one type of an enantiomeric pair is more lost by collisions with particles with weak charge in space or on the earth [6]. Another example is the Chirality-Induced Spin Selectivity (CISS) [13]. CISS is the strong spin polarization when electrons pass through chiral organic molecules. It has not been clarified what makes large spin polarization. For this puzzle, the study from the viewpoint of the torque may be a key to the solution instead of the spin-orbit interaction, which is in the concept of energy. The torque for the electron spin is dependent on the electron chirality density, and the torques induced by the left- and right-handed electrons are opposite. ECCM predicts opposite electron chirality distribution in enantiomeric pair molecules. The zeta force may be the driving torque for the CISS. The correlation between the spin polarization of the CISS and the magnitude of the zeta force was confirmed for alanine and helicene molecules [14].
@@
In the present paper, it is confirmed that the enhancement by electronic excitation also occurs for the MEC. As target molecules, H2Te2 and H2Se2 are chosen. In Refs. [7, 8], H2O2 and H2S2 are also studied for the enhancement of the PVED. For the MEC, the computation of molecules including only light elements with small spin-orbit interaction such as H2O2 and H2S2 requires very large basis sets and high cost [6, 11]. Hence, H2O2 and H2S2 are omitted in the present study for the excited states. However, we consider that H2O2 and H2S2 will show results similar to H2Te2 and H2Se2 due to the similar patterns of the total chirality in the ground state [6, 11] and the enhancement of the PVED [7, 8]. After the enhancement of the MEC is confirmed, it is studied whether the CBE hypothesis can also be applicable to the enhancement of the MEC.
@@ II Theory
Here $E_{\rm PV,HOMO}$ is the contribution from one electron in the Kramers pair of the HOMO to $E_{\rm PV}$. (Two electrons in the same Kramers pair give the same contribution to the electron chirality.) As a more improved estimate, contributions from other orbitals are taken into account,
$E_{\rm PV}({\rm IE})=\sum_{i}^{\rm unocc}r_{i}^{U}E_{\rm PV,\it i}-\sum_{i}^{\rm occ}r_{i}^{O}E_{\rm PV,\it i},$ (7)
where $E_{\rm PV,\it i}$ is the contribution from one electron in the $i$-th occupied or unoccupied orbital, and $r_{i}^{O}$ ($r_{i}^{U}$) is the excitation ratio of the electron in the $i$-th occupied orbital (to the $i$-th unoccupied orbital). The former simple estimate is quite useful, since only Hartree-Fock (HF) or density functional theory computations are needed, while the improved estimate (IE) requires a computation beyond self-consistent field computations.
@@
where $\psi_{\rm HOMO}$ is the wave function of the HOMO. The improved estimate with other orbital contributions is given in the following expression,
$\sum_{i}^{\rm unocc}r_{i}^{U}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x)-\sum_{i}^{\rm occ}r_{i}^{O}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x),$ (9)
where $\psi_{i}$ is the wave function of one of the Kramers pair of the $i$-th orbital. In this paper, the CBE hypothesis is checked for the MEC.
@@ IV Results
Figure 1 shows the electron chirality in excited states of (a) H2Te2 and (b) H2Se2 as a function of the dihedral angle. The notation of excited states, $nx$, means the $n$-th excited state obeying the $x$-symmetry. The electron chirality in excited states have opposite sign and much larger values compared to that of the ground state (GS). The enhancement of the MEC as well as the PVED is confirmed. While the MEC in the ground state on the dihedral angle is a sigmoidal curve, in excited states a different pattern of the dependence is seen. All curves of the MEC have similar distributions. In their patterns, the MEC increases to 90 and flips the sign around $\phi=90^{\circ}$. Most excited states are derived as the excitation of one electron in the HOMO in the ground state. The excitation ratio of the HOMO is roughly 90% except for 2a and 2b states around $\phi=90^{\circ}$ for both molecules, where the ratio is 20-70%. The energies of these excited states were reported in Ref. [8]. For both molecules, 1a, 1b, and 2a ($\phi<90^{\circ}$) or 1a, 1b, and 2b ($\phi>90^{\circ}$) are in the same approximate triplet state. The rest 2a (2b) state is the singlet state for $\phi>105^{\circ}$ $(\phi<75^{\circ})$, and forms approximate triplet state with 3a and 3b for $90^{\circ}<\phi<105^{\circ}$ $(75^{\circ}<\phi<90^{\circ})$.
@@
In Fig. 3, the estimate of the MEC in the CBE hypothesis is shown as well as the improved estimate as introduced in Sec. II. The predictions by the CBE hypothesis well realize the patterns of the MEC in 1a and 1b states. Particularly, all predictions for $\phi>90^{\circ}$ are perfectly consistent with computational results of the MEC. For $\phi<90^{\circ}$ the prediction in the 1a state of H2Te2 realizes precisely the MEC curve, while prediction curves for other states have a few hundred percents deviations. This inconsistency is considered to arise from the modification of occupied orbitals from the ground state and the difference between HF and CCSD computations. Nevertheless, the patterns of curves are consistent with the computational results. The CBE hypothesis is considered to apply the MEC. At most points, the simple CBE estimate does not deviate from the prediction by IE. This is because the electron in the HOMO is dominantly excited at most points. The prediction by the improved estimate seems to be worse than the simple CBE estimate around $\phi=90^{\circ}$. This is because there are few data for IE points. If there were additional data around $\phi=90^{\circ}$, the IE prediction was better than Fig. 3.

### v2

@@ I Introduction
It is well established that certain molecules exhibit chiral structures, while electrons also possess an intrinsic chirality, i.e. the handedness of the electron can be defined. A free electron does not have a specific charge distribution and the handedness of the electron should not be defined by the distribution. However, the electron possess spin, and the inner product of spin and momentum enables the definition of a specific handedness quantity known as helicity. In the absence of external fields, helicity is conserved over time, since it commutes with the Hamiltonian operator. It is noteworthy that helicity is frame-dependent and is not Lorentz invariant. In contrast, the chirality of an electron is defined as a quantity that maintains a consistent handedness ratio across all inertial frames. This concept represents a relativistic extension of helicity as elucidated in the next section.
@@
Nature inherently selects the chirality as the handedness quantity of electrons. Notably, the weak interaction, one of the four fundamental interactions in particle physics, differentiates between left- and right-handed electrons based on their chirality. For instance, this interaction is mediated by W and Z gauge bosons, and the W boson interacts with left-handed electrons but does not with right-handed electrons. Additionally, in the framework of quantum field theory, the torque density for the electron spin is predicted to be dependent on the electron chirality density [1, 2]. In this theoretical prediction, the local torque contribution, which is named the zeta force, is expressed as the gradient of the electron chirality density. While this contribution to torque density has yet to be observed experimentally, it has been corroborated by numerical simulations [3].
@@
Although the chirality of an electron affects its physical properties, free electrons eventually lose their chirality polarization due to mass-induced chirality flips. For instance, if electrons are initially polarized predominantly in the left-handed chirality, this polarization will decay over time to a state of no net polarization. The left-handed component tends to convert to the right-handed component more frequently than the reverse process occurs, as the initial ratio of left-handed to right-handed components is greater.
@@
However, electrons within chiral molecules, which are parity-violating entities, are expected to exhibit polarization in chirality. This is because a state polarized in electron chirality may represent the most stable configuration in steady-state conditions due to spin-orbit interactions. Although the mass of the electron can induce chirality flips, the ultimate objective is to achieve a stable polarized state in chirality.
@@
One phenomenon associated with the electron chirality is the parity-violating energy difference (PVED) between enantiomers. This energy difference arises from the Z-boson exchange interaction between nuclei and electrons and is proportional to the difference in electron chirality density at the nuclei of the enantiomers. Although extensive experimental efforts have yet to confirm this energy difference [4], theoretical studies utilizing quantum chemistry computations indicate a nonzero PVED in many chiral molecules [5, 6, 7, 8]. Importantly, this energy difference pertains solely to the electron chirality at the nuclei. The electron chirality for the entire molecule of H2Te2 was reported to be nonzero [9]. Furthermore, systematic analyses in Ref. [6, 10, 11] demonstrate that chiral molecules generally exhibit a nonzero integrated value of electron chirality density over the entire molecular structure. The existence of this chirality is named Electron Chirality in Chiral Molecules (ECCM) [11].
@@
The value of ECCM is small, and phenomena related to ECCM are not extensively documented. Nevertheless, our research group has investigated several associated phenomena. One topic is the homochirality in nature [12]. In an enantiomeric pair, the number of the left-handed electron differs, while the total number of electrons is the same. Consequently, one enantiomer exhibits a higher reaction rate for weak interactions than its counterpart. This disparity leads to the preferential loss of one enantiomer through collisions with weakly charged particles in both terrestrial and extraterrestrial environments [6]. Another noteworthy phenomenon is Chirality-Induced Spin Selectivity (CISS) [13]. CISS is the strong spin polarization when electrons traverse chiral organic molecules such as DNA. The mechanisms underlying this significant spin polarization remain unclear. For this puzzle, examining the problem from the perspective of torque may provide critical insights, as opposed to focusing solely on spin-orbit interactions, which are typically framed in terms of energy. The torque experienced by electron spin is influenced by electron chirality density, with the torques generated by left- and right-handed electrons being oppositional. ECCM predicts an opposite distribution of electron chirality within enantiomeric pair molecules, suggesting that the zeta force may serve as the driving torque for CISS. The correlation between spin polarization in CISS and the magnitude of the zeta force was confirmed for alanine and helicene molecules [14].
@@
In the present paper, it is confirmed that electronic excitation also leads to an enhancement of MEC. Target molecules chosen for this investigation are H2Te2 and H2Se2. These H2X2 type molecules are often employed to study parity-violating effects in molecules [9, 17, 20]. In Refs. [7, 8], H2O2 and H2S2 are also studied for the enhancement of the PVED. However, computing MEC for molecules that consist solely of light elements with small spin-orbit interactions, such as H2O2 and H2S2 necessitates large basis sets and incurs significant computational costs [6, 11]. Specifically, in Ref. [6], it was demonstrated that the dyall.ae3z basis set is insufficient for accurately calculating the MEC of H2S2, while the dyall.ae4z basis set is required for reliable results. Moreover, for the calculation of the MEC of H2O2, even the dyall.ae4z basis set is insufficient. The accurate calculation of H2O2 requires additional diffuse functions [11]. Consequently, H2O2 and H2S2 are excluded from the present study focused on excited states. Nonetheless, we anticipate that H2O2 and H2S2 will exhibit results analogous to those of H2Te2 and H2Se2 due to the similar patterns of the total chirality in the ground state [6, 11] and the enhancement of the PVED [7, 8]. Following the confirmation of MEC enhancement, it is studied whether the CBE hypothesis is applicable to the enhancement of the MEC.
@@ II Theory
Here $E_{\rm PV,HOMO}$ is the contribution from one electron in the Kramers pair of the HOMO to $E_{\rm PV}$. (Two electrons in the same Kramers pair give the same contribution to the electron chirality.)
@@
where $\psi_{\rm HOMO}$ is the wave function of the HOMO. In this paper, the CBE hypothesis is checked for the MEC.
@@ IV Results
Figure 1 shows the electron chirality in excited states of (a) H2Te2 and (b) H2Se2 as a function of the dihedral angle. The notation of excited states, $nx$, means the $n$-th excited state obeying the $x$-symmetry. The electron chirality in excited states have opposite sign and much larger values compared to that of the ground state (GS). The enhancement of the MEC as well as the PVED is confirmed. While the MEC dependence on the dihedral angle in the ground state is a sigmoidal curve, in excited states a different pattern of the dependence is seen. All curves of the MEC have similar distributions. In their patterns, the MEC increases to 90 and flips the sign around $\phi=90^{\circ}$. The structure at $\phi=90^{\circ}$ is chiral, while the MEC is zero for the GS and some excited states. The reason why the PVED is zero around $\phi=90^{\circ}$ in the ground state is explained in terms of the parity violation operator $\sigma\cdot p$ and spin-orbit interaction using perturbation theory in Ref. [33]. Since the MEC and PVED has the same dependence on $\phi$, the same discussion will clarify why the MEC is zero around $\phi=90^{\circ}$ in the ground state. Most excited states are derived as the excitation of one electron in the HOMO in the ground state. The excitation ratio of the HOMO is roughly 90% except for 2a and 2b states around $\phi=90^{\circ}$ for both molecules, where the ratio is 20-70%. The energies of these excited states were reported in Ref. [8]. For both molecules, 1a, 1b, and 2a ($\phi<90^{\circ}$) or 1a, 1b, and 2b ($\phi>90^{\circ}$) are in the same approximate triplet state. The rest 2a (2b) state is the singlet state for $\phi>105^{\circ}$ $(\phi<75^{\circ})$, and forms approximate triplet state with 3a and 3b for $90^{\circ}<\phi<105^{\circ}$ $(75^{\circ}<\phi<90^{\circ})$.
@@
In Fig. 3, the estimate of the MEC in the CBE hypothesis is shown. The predictions by the CBE hypothesis well realize the patterns of the MEC in 1a and 1b states. Particularly, all predictions for $\phi>90^{\circ}$ are perfectly consistent with computational results of the MEC. For $\phi<90^{\circ}$ the prediction in the 1a state of H2Te2 realizes precisely the MEC curve, while prediction curves for other states have a few hundred percents deviations. This inconsistency is considered to arise from the modification of occupied orbitals from the ground state and the difference between HF and CCSD computations. Nevertheless, the patterns of curves are consistent with the computational results. The CBE hypothesis is considered to apply the MEC.